# American Option Pricing Under Time-Varying Rough Volatility: A Signature-Based Hybrid Framework

### v1

@@ 4 Example Empirical Results and Model Behavior
We tested the hybrid architecture on many options contracts. Data was sourced from OptionMetrics via WRDS [19]. Two representative examples selected were two 10-day American put options (AAPL and META) from August 31, 2023.
@@ 5 Code Implementation
To support reproducibility and future extensions, the complete implementation of the proposed framework is available on GitHub at roshanshah11/American-Option-Pricing. The repository includes source code for volatility estimation, Hurst forecasting, simulation under multiple engines, signature extraction, and primal–dual valuation. A detailed walkthrough of the architecture and numerical experiments is provided in the accompanying Jupyter notebook, available at Notebooks/final.ipynb. The framework operates on two primary datasets: one consisting of listed options contract data with fields such as date, days, forward_price, strike_price, premium, impl_volatility, cp_flag, ticker, and index_flag [19]; and a second dataset containing historical equity prices in the format (date, ticker, close, return) [20].
Acknowledgements
Thank you to Ms. Patel for her valuable guidance and feedback throughout the research and writing process. This work also benefited from conversations with faculty at The Lawrenceville School and from access to OptionMetrics data via the Wharton Research Data Services (WRDS) platform. All errors are the author’s own.

### v2

@@ 4 Example Empirical Results and Model Behavior
We tested the hybrid architecture on many options contracts. Data was sourced from OptionMetrics via WRDS (OptionMetrics 2025a). Two representative examples selected were two 10-day American put options (AAPL and META) from August 31, 2023.
@@ 5 Code Implementation
To support reproducibility and future extensions, the complete implementation of the proposed framework is available on GitHub at https://github.com/roshanshah11/American-Option-Pricing-Under-Time-Varying-Rough-Volatility. The repository includes source code for volatility estimation, Hurst forecasting, simulation under multiple engines, signature extraction, and primal–dual valuation. A detailed walkthrough of the architecture and numerical experiments is provided in the accompanying Jupyter notebook, available at https://github.com/roshanshah11/American-Option-Pricing-Under-Time-Varying-Rough-Volatility/blob/main/Notebooks/final.ipynb. The framework operates on two primary datasets: one consisting of listed options contract data with fields such as date, days, forward_price, strike_price, premium, impl_volatility, cp_flag, ticker, and index_flag (OptionMetrics 2025a); and a second dataset containing historical equity prices in the format (date, ticker, close, return) (OptionMetrics 2025b).
Acknowledgments
Thank you to Ms. Pooja Patel for her valuable guidance and feedback throughout the research and writing process. This work also benefited from conversations with faculty at The Lawrenceville School and from access to OptionMetrics data via the Wharton Research Data Services (WRDS) platform. All errors are the author’s own.